OpenStudy (anonymous):

Find the product. (-5a^2)^3·a^5

OpenStudy (whpalmer4):

Here you get another property of exponents: \[(ab)^n = a^nb^n\] So you can rewrite the first part as \[(-5a^2)^3 = (-5)^3(a^2)^3\] and use what you learned on the previous problem to do the rest.

OpenStudy (anonymous):

I tried solving it and my answer didn't fit with the answers provided... That's why I posted it. x) I'm not very good at this. Hoping someone could help me with the answer. Maybe walk thru how to do it?

OpenStudy (whpalmer4):

Okay, that's what I'm trying to do, but perhaps the math notation isn't clear to you. Continuing on from where I left off, can you figure out the value of \((-5)^3\)?

OpenStudy (whpalmer4):

Hint: it's the same as (-5)(-5)(-5)

OpenStudy (anonymous):

So, -125?

OpenStudy (whpalmer4):

Very good. So now we have \[(-5a^2)^3 *a^5 = -125 (a^2)^3 a^5\]Can you simplify \((a^2)^3\)?

OpenStudy (whpalmer4):

Hint: \[(a^n)^m = a^{n*m}\]

OpenStudy (anonymous):

a^6?

OpenStudy (whpalmer4):

Yes! If that doesn't make complete sense to you, look at it this way: \[(a^2)^3 = (a*a)^3 = (a*a)*(a*a)*(a*a) = a*a*a*a*a*a = a^6\]

OpenStudy (whpalmer4):

Now we've got \((-5a^2)^3*a^5 = -125a^6a^5\) Can you finish the simplification?

OpenStudy (whpalmer4):

Here the visual model would be \[a^6a^5 = (a*a*a*a*a*a)*(a*a*a*a*a)\]\[ = a*a*a*a*a*a*a*a*a*a*a = a^?\]

OpenStudy (anonymous):

-125a^11???

OpenStudy (whpalmer4):

That's the final answer, yep!

OpenStudy (anonymous):

Thank you so much!!! :D

OpenStudy (whpalmer4):

They aren't really that hard, it just takes some practice (and attention to detail). Here's one for you, let's see if you can build on what you've learned: What is \(((a^2)^3)^4\)?

OpenStudy (anonymous):

a^24?

OpenStudy (whpalmer4):

Ding ding ding! We have a winner!

OpenStudy (anonymous):

Yay! I wish you were my actual teacher. She left us with a sub that doesn't know like... Anything! And my teacher isn't much help. She's just like, "Here's the problem, here's the answer."

OpenStudy (anonymous):

You helped me understand!

OpenStudy (whpalmer4):

Great! That's always my goal :-) It's nice working with people who want to understand, not just have an answer to put into their computer...

OpenStudy (anonymous):

Thank you @whpalmer4. You're such a great teacher!

OpenStudy (whpalmer4):

Thanks, you made my day!

OpenStudy (anonymous):

I'm glad! :D